OpenStudy (anonymous):

How can I write an equation of a line in standard form given the slope and y int.? :)

OpenStudy (anonymous):

Is there a formula to follow?

OpenStudy (anonymous):

y = mx + b :) Just substitute slope m and y - intercept b and ur answer is framed :)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

@ganibl Wait a sec. though, my answer has to look like Ax+By=C

OpenStudy (anonymous):

Whats the slope and y -intercept given to U??

OpenStudy (anonymous):

y + mx = b :) It would be of the form Ax + By = C :)

OpenStudy (anonymous):

slope -1/7 y int. 0.2

OpenStudy (anonymous):

Yeah I just got it

OpenStudy (anonymous):

y = mx + b y = -1/7x + 0.2 7y = -x + 0.2 7y + x = 0.2 x + 7y = 0.2

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

Thanks again:)